Lynn invests 10,000$. in one account, he earns 6% interest, in the other, he earns 4.5%. How much did he invest in each account if he makes 570$ in one year?
PLEASE HELP!!!!!!!!!!!

Answer by checkley71(8403)

.06X+.045(10000-X)=570
.06X+450-.045X=570
.015X=570-450
.015X=120
X=120/.015
X=8000 INVESTED @ 6%
10000-8000=2000 INVESTED @4.5%
PROOF
8000*.06+2000*.045=570
480+90=570
570=570